Solve the following linear equations. 
(x - 5)/3 = (x - 3)/5 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(x-5)/3=(x-3)/5
⇒ 5(x-5) = 3(x-3)
⇒ 5x-25 = 3x-9
⇒5x-3x=-9+25
⇒2x=16
⇒x=8

Thanks